OpenStudy (andriod09):

Math help to test out of Alg 2, review work: equation in the comments!

OpenStudy (andriod09):

\[(y - 1)(y^2 + 2y + 1)\]

OpenStudy (anonymous):

what are we trying to do here?

OpenStudy (andriod09):

Have to solve the polynomal by distributing.

OpenStudy (jhannybean):

Use the method of grouping. \[\large (y)(y^2+2y+1) -(1)(y^2+2y+1)\]

OpenStudy (jhannybean):

\[\large y^3 +2y^2 + y -y^2 -2y -1\] Can you simplify that?

OpenStudy (andriod09):

Question, if you have \[(y)(y^2)\] wouldn't that be \[2y^2\]?

OpenStudy (jhannybean):

Nope.

OpenStudy (jhannybean):

\[\large y\cdot y^3 = (1\cdot 1) \cdot y^ {1+2}\]

OpenStudy (jhannybean):

@Loser66 I guess distributing just helps break down the equation into simplified components. Good enough to evaluate.

OpenStudy (jhannybean):

I don't know if we do have to factor again, he just asked to simplify the equation?

OpenStudy (loser66):

@Jhannybean read the 3rd comment :"I have to solve...."

OpenStudy (jhannybean):

@andriod09 \[\large 2y^2 = y\cdot 2y = (2\cdot 1)\cdot y^{1+1}\]

OpenStudy (andriod09):

I have to put it into lowest terms. yes, i saw it. I'm letting you two finish your discussion. >.<

OpenStudy (jhannybean):

And @Loser66 I'm just doing what i think is the easiest method :P

OpenStudy (jhannybean):

@completeidiot

OpenStudy (loser66):

@Jhannybean we have to know what we are supposed to do first, right?

OpenStudy (anonymous):

you cant solve for x because its not an equation

OpenStudy (jhannybean):

How would you go about solving this? @Loser66

OpenStudy (anonymous):

i would agree with jhanny's method to just expand the expression because question did not say to factor

OpenStudy (loser66):

because the asker said at the 3rd comment:"I have to solve...", re read, please

OpenStudy (andriod09):

@Loser66 I clarified that by solving, i had to simplify it.

OpenStudy (jhannybean):

Solving doesn't necessarily mean solving for a variable.... it might as well mean factor it to the simplest form.

OpenStudy (jhannybean):

Shall we continue, @andriod09

OpenStudy (andriod09):

Its a polynomal, which i just recently figured out, as I got it wrong when i solved for x.

OpenStudy (andriod09):

Yes, please.

OpenStudy (jhannybean):

And so to clarify this,we're just trying to simplify this problem into the most basic form, yes?

OpenStudy (andriod09):

Yes, we're simplifying.

OpenStudy (jhannybean):

so what do you get when you simplify this? \[\large y^3 +2y^2 + y -y^2 -2y -1\]

OpenStudy (andriod09):

OKay, give me a minute, i have got to switch brosers, this one is lagging hard. Brb.

OpenStudy (andriod09):

BAck. Wouldn't you get: \[2y^3-1\]? i added up the \[y^3 + 2y^2 + y\] which gave me: \[4y^5\] then i subtracted \[3y^2\] Is that right?

OpenStudy (jhannybean):

No. These are all different like terms...

OpenStudy (andriod09):

What do you mean?

OpenStudy (jhannybean):

Just remember - the only timewe an add powersis when we're multiplying 2 terms. ex: \(\large y \cdot y^3 = y^{3+1}\)

OpenStudy (jhannybean):

time we can add*

OpenStudy (jhannybean):

So here, you would have \[\large y^3 +2y^2 + y -y^2 -2y -1\]\[\large y^3 +(2y^2 -y^2)+(-2y +y) -1\]

OpenStudy (andriod09):

So would That be the answer?

OpenStudy (jhannybean):

Simplify it.

OpenStudy (andriod09):

How?

OpenStudy (anonymous):

adding and subtracting like terms

OpenStudy (jhannybean):

\[\large y^3 +(2y^2 -y^2)+(-2y +y) -1\]\[\large (1)y^3 +y^2(2-1)+y(-2+1)-1\] Now you can reduce that i'm sure.

OpenStudy (andriod09):

would it be: \[y^3 + 2y^2 - y^2 - 2y + y -1\]?

OpenStudy (jhannybean):

you can reduce 2y^2 -y^2 and -2y +y like i showed you above ^^

OpenStudy (andriod09):

If you haven't I REALLY hate polynomals. I don't understand them.

OpenStudy (jhannybean):

Just evaluate this.....\[\large (1)y^3 +y^2(2-1)+y(-2+1)-1\] you'll be done.

OpenStudy (andriod09):

I don't get it still. What do i do to it? because every time i try to do something, its wrong.

OpenStudy (jhannybean):

Are yu getting \[\large y^3 +y^2 -y-1\]

OpenStudy (andriod09):

No. I"m not. I'm getting what I told you.

OpenStudy (jhannybean):

Ok... \[\large y^3 + \color{red}{2y^2 - y^2}\color{blue}{ - 2y + y} -1\] you can simplify these because these two highled portions are "like-terms" they both have the same power.

OpenStudy (andriod09):

Oh, so they would cancel each other? I "knew" that. >.<

OpenStudy (jhannybean):

\[\large 2y^2 -y^2\] are BOTH to the power of 2, and because these polynomials are ADDING eachother they can be simplified.

OpenStudy (andriod09):

Ohhhhhhh. I'm stupid.

OpenStudy (jhannybean):

Nah,Probably just a brain fart.

OpenStudy (jhannybean):

Do you get how to reduce it now?

OpenStudy (andriod09):

Yes. Time to go be a jerk to my other problems. I'll be back if i need more help

OpenStudy (jhannybean):

Sure thing :)

OpenStudy (andriod09):

So wait, what is the answer? My STM just kicked in.

OpenStudy (jhannybean):

STM?

OpenStudy (jhannybean):

\[\large y^3 +y^2 -y-1\] would be the most simplified form.

OpenStudy (andriod09):

Ah okay. Thanks.

OpenStudy (andriod09):

I'l pm you what STM means.